OpenStudy (anonymous):

Solve for x: (x - 9)2 = 1

OpenStudy (anonymous):

i assume you mean (x-9)^2=1 then we have x-9=1 or x-9=-1 by taking the square root of both sides solving both equations, we have x=10 or x=8

OpenStudy (anonymous):

x-9=1/2 x=1/2+9 x = 11/9

OpenStudy (anonymous):

thank you marc